CLINICAL CHEMISTRY:

Assuming carbon dioxide is kept at a constant partial pressure of 55 mmHg, and your starting bicarbonate concentration is 22 mM; calculate how much excess bicarbonate it would require in order to raise the pH by .05 units.
Answer #1

We know the relation between pH, conc of bicarbonate and parital pressure of CO2 is

pH = 6.1 + log [HCO3-] / 0.03pCO2

Given :

[HCO3-] = 22mM

pCO2 = 55mmHg

pH = 6.1 + log 22/ 0.03 X 55 = 7.22

Now we want to raise the pH by 0.05 units

new pH = 7.22 + 0.05 = 6.1 + log [HCO3-] / 0.03X pCO2

7.27 -6.1 = log [HCO3-] / 0.03 x pCO2

14.79 = [HCO3-] / 0.03 X pCO2

[HCO3-] = 24.40 mM

so excess bicarbonate needed = 24.40 - 22 mM = 2.4mM